If a=(1,-1,1) and c=-1,-1,0) then the vector b satisfying a*b=c and a.b=1 

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear  palllavi

 a = i -j + k

c = -i -j

aXb=c

or

 a X (a X b) = a X c

 (a.b) a  - (a.a) b = a Xc

 a - 3b  = i -j -2k

  b = k

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin